OpenStudy (anonymous):

Find an equation of the plane. The plane passes through the points (5, 3, 1) and (5, 1, -5) and is perpendicular to the plane 8x + 9y + 2z = 11.

OpenStudy (anonymous):

I found the vector of two points, crossed points w/ a,b,c from eq of plane. then I plugged into eq. of plane formula. what did I do wrong?

OpenStudy (anonymous):

Check your answer by taking dot product of two perpendicular vectors.

OpenStudy (anonymous):

I see a mistake in my calculation.

OpenStudy (anonymous):

What is wrong in my method?

OpenStudy (anonymous):

Show us your method.

OpenStudy (anonymous):

(5,3,1) (5,1,-5) -> <0,-2,-6> <0,-2,-6>X<8,9,2>=<50,-48,16> <50,-48,16>.<8,9,2> = 0 <50,-48,16>=<25,-24,8> 25(x-5)^2-24(y-3)^2+8(z-1)=0 25x-125-24y-72+8z-8=0 25x-24y+8z=205

OpenStudy (anonymous):

What does your answer say in book, apparently you are checking it?

myininaya (myininaya):

chag (sorry to interrupt) did you ask about the calculus dream problem or was that someone else

OpenStudy (anonymous):

Someone else.

myininaya (myininaya):

k thanks

OpenStudy (anonymous):

There is another method, I don't know if this is valid <a,b,c>dot V=0. it gives you 25x-27y+2z=0

OpenStudy (anonymous):

Not correct

OpenStudy (anonymous):

You asked what is wrong with your method. The mistake is when you did the equation formula you squared x and y; you shouldn't have squared it 25x-24y+8z=? Plug in (5,3,1) and you get 25x-24y+8z=61